OpenStudy (anonymous):

The formula can be used to find the volume of a rectangular pyramid whose base area is b square units and whose height is h units. What is the height of a pyramid whose volume is 360 cm3 and whose base area is 72 cm2? _________ cm

OpenStudy (sleepyjess):

Do you know the formula for volume of a pyramid?

OpenStudy (anonymous):

V=lwh/3

OpenStudy (sleepyjess):

Yep, so we can substitute what we already know into that.

OpenStudy (sleepyjess):

l*w is equal to the base area.

OpenStudy (sleepyjess):

So now we have \(\bf 360=\dfrac{72*h}{3}\)

OpenStudy (sleepyjess):

What do we do to get rid of the 3?

OpenStudy (anonymous):

no idea

OpenStudy (sleepyjess):

Multiply both sides by 3.

OpenStudy (sleepyjess):

What is 360*3?

OpenStudy (anonymous):

1080

OpenStudy (sleepyjess):

Yes, now we have \(\bf 1080=72*h\). What do you think we do to get rid of 72?

OpenStudy (sleepyjess):

Remember, use the opposite operation of multiplying.

OpenStudy (anonymous):

divide it by 3?

OpenStudy (sleepyjess):

3 is gone now.

OpenStudy (sleepyjess):

You have the right idea, divide, but the incorrect number.

OpenStudy (anonymous):

divide by 72?

OpenStudy (sleepyjess):

Yep

OpenStudy (sleepyjess):

What do you get when you divide by 72?

OpenStudy (anonymous):

0

OpenStudy (sleepyjess):